A firm is faced with undertaking one of these four projects in 2022 – open new branch, introduce another product, purchase a brand new vehicle, or increase workers’ salary. The costs and expected benefits of the projects are presented in the table below.

  Activity

Branch Product Vehicle Salary

Direct cost Benefit (GH¢) (GH¢) 380 390 400 400 350 392 422 425

        Use the information provided in the table to answer the following questions.

a) Compute the net benefit of each project using the direct cost.

b) Find the opportunity cost of each project.

c) Compute the net benefit of the projects using the opportunity cost.



1
Expert's answer
2022-03-17T14:12:03-0400

Question 1

The benefit using the direct cost

The net benefit for each project is the Benefit - the direct cost.

Branch

350-380= -30

Product

392-390= 2

Vehicle

422-400= 22

Salary

425-400= 25


Question 2

The opportunity cost of each project

  • For Branch

The opportunity cost = Return on the other projects- Return on the chosen project

For Branch

= (2+22+25)-(-30)= 79


  • For product

=(-30+22+25)-2 = 15


  • For Vehicle

(-30+2+25)-22= -25


  • For salary

(-30+2+22)-25= -31


Question 3

The net benefit using the opportunity cost.

For Branch

350--79= 429

For Product

392-15= 377

For Vehicle

422--25 = 447

For Salary

425--31= 456